A student club consists of 14 girls and 16 guys. A committee of 8 is to be selected. In how many ways can this be done if: A.) the committee must have at least 6 girls on it. B.) three officers (president, vice president, secretary/treasurer) are selected from ALL the members. Please show work.
Answer #1

Ans:

A)

Case 1)When 6 girls,2 guys

Number of ways =14C6*16C2=360360

Case 2)When 7 girls,1 guy

Number of ways =14C7*16C1=54912

Case 2)When 8 girls

Number of ways =14C8=3003

Total number of ways=360360+54912+3003=418275

B)In this case order matters,so we use permutation.

Number of ways=30P3=24360